NEXT
Make New Entry,
Make Followup Entry
User name riordan
Log entry time 17:46:34 on March 01, 2009
Entry number 264097
This entry is a followup to: 264093
keyword=Summary for Feb 28 Swing - March 1 Day
Summary for Feb 28 swing - Mar 1 Day
Finishing 5-pass production on Kin I. Starting Kin G.
Beam quality has been poor preventing smooth data taking. For the
remaining kinmatics, we anticipate only getting about 50% of the
orginially planned data.
Beam in Use
shift |
ABU |
BANU |
BNA |
2009/02/28 swing | 4.5 |
1.0 | 2.5 |
2009/03/01 owl | 2.5 |
0.6 | 4.9 |
2009/03/01 day | 4.0 |
1.1 | 2.9 |
total | 11.0 |
2.7 | 10.3 |
Main events/issues:
- Feb 28
- Production on 5-pass Kin I, p0 = 1.34 GeV/c
- Beam quality is very poor due to RF problems, injector
problems, and vaccuum problems.
- Mar 1
- Brad and Kalyan opportunistically added preshower to T6
trigger for explicit pion rejection, increasing data taking rates.
- Had left Q1 trip at 8:00 due to low liquid levels. Brad
noticed that the readback levels were fluctuating significantly.
- No Compton data from 7:00~13:00 due to Compton DAQ problems.
- 14:50 - Switching to Kin G, p0 = 1.2 GeV/c
Plans until Beam Studies, 3/3
- Continue production G, E until Monday, 3/2 6:30am
- 6:30, go to controlled access for techs to swich LHRS polarity
- 9:00-17:00 - 5 pass positron runs
- 19:00 - Moller
- Finish as much 5 pass production as possible until beam studies
Tuesday in Kin E, C
A copy of this log entry has been emailed to: brads, jpchen, jiang, meziani, choi